OpenStudy (anonymous):

Which expressions show how 8 • 54 can be rewritten using the distributive property? Choose all answers that are correct. A. 8 • 50 + 4 B. 8 • 50 + 8 • 4 C. 8 • 60 – 8 • 6 D. 8 • 60 + 8 • 4

OpenStudy (anonymous):

B because: \[8\times(54)=8\times(5+4)=8\times(5)+8\times(4)\]

OpenStudy (anonymous):

ty 8)